### v1

# Robust Latent Representation Tuning for Image-text Classification

Abstract
Large models have demonstrated exceptional generalization capabilities in computer vision and natural language processing. Recent efforts have focused on enhancing these models with multimodal processing abilities. However, addressing the challenges posed by scenarios where one modality is absent remains a significant hurdle. In response to this issue, we propose a robust latent representation tuning method for large models. Specifically, our approach introduces a modality latent translation module to maximize the correlation between modalities. Following this, a newly designed fusion module is employed to facilitate information interaction between the modalities. In this framework, not only are common semantics refined during training, but the method also yields robust representations in the absence of one modality. Importantly, our method maintains the frozen state of the image and text foundation models to preserve their abilities acquired through large-scale pretraining. We conduct experiments on several public datasets, and the results underscore the effectiveness of our proposed method.
Index Terms—  Image-text classification, large models, robust learning, representation learning.
1 Introduction
In recent times, large models have garnered substantial attention due to their remarkable generalization capabilities across numerous downstream tasks. Given that most large models are pretrained on unimodal datasets (e.g., LLaMA [1], OPT [2]), researchers have sought to augment these models with multimodal processing capabilities. Notably, approaches like LISA [3] have proposed extracting multimodal features using various large models, employing these features for tasks such as image segmentation. PixelLM [4] has introduced a tuning framework wherein visual embeddings are prefixed to textual tokens, jointly processed by large language models (LLM). Despite the numerous endeavors to imbue large models with the capacity to process multimodal signals (e.g., images and texts), there has been limited attention to robust representation learning, and performance in modality-absence scenarios remains relatively unexplored. Real-world applications frequently encounter scenarios where certain modalities are absent, rendering current methods challenging to apply.
To address this challenge, we introduce a novel strategy for robust multimodal representation tuning in this paper. Our approach leverages two pretrained large models dedicated to image and text processing. At each corresponding layer of the paired image-text models, we incorporate a Modality Latent Translation (MoLT) module. Within this module, image and text embeddings are projected onto a shared latent space, aiming to bring the embeddings closer together. This shared space acts as a bridge connecting the image and text domains. Subsequently, a cross-attention mechanism is employed after feature extraction to capture the relationship between the robust representation and the associated modality embeddings for making predictions.
At the heart of our method, MoLT comprises two cross-attention modules, individually tailored for the image and text domains. Following the cross-attention step, we apply a Canonical Correlation Analysis (CCA) loss [5] to facilitate the learning of a robust representation between the two modalities. Consequently, in scenarios where one modality is absent, a straightforward translation from the available modality or the utilization of the learned robust representation becomes feasible for downstream tasks. Throughout our training process, the parameters from pretrained models remain frozen, allowing only the newly introduced modules to be tunable. This approach enables the model to progressively acquire and refine robust representations.
In summary, our contributions can be outlined as follows:
  • We propose a novel strategy for robust representation tuning in large models. Our method facilitates the learning of a robust representation in a shared latent space, establishing a bridge between image and text embeddings.
  • Introducing the Modality Latent Translation (MoLT) module in our approach, we present a sophisticated cross-attention module that brings text and image embeddings closer together.
  • Our model achieves state-of-the-art performance on evaluated image-text classification datasets. Furthermore, our experiments demonstrate the model’s remarkable robustness in scenarios involving modality absence.
2 Related Prior Works
2.1 Large Vision and Language Models
The advent of large models has dominated discussions in deep learning, particularly within the realms of computer vision and natural language processing. Noteworthy language models include GPT-3 [6], LLaMA [1], and OPT [2], which, pretrained on extensive corpora, exhibit formidable capabilities in comprehending and generating long-context information. In the domain of computer vision, SAM [7] stands as the current state-of-the-art foundation model for visual understanding. However, the scarcity of open-source large models trained on multimodal corpora, such as CLIP [8], poses challenges for processing multimodal data using large models.
2.2 Multimodal Large Model Tuning
Recent years have witnessed a surge of interest in the tuning of large models. While most tuning strategies are devised for unimodal processing, some researchers have endeavored to integrate multimodal information into large models through multimodal tuning. For instance, Flamingo [9] proposes fusing multimodal signals with gated cross-attention into a frozen image encoder, showcasing the potential of large models for multimodal processing. BLIP [10] aligns multimodal embeddings through multitask learning, and BLIP-2 [11], subsequently proposed with a Q-Former, finds widespread application in recent works. PaLM-E [12] introduces sending visual tokens as input to pretrained language models, demonstrating impressive performance. In FROMAGe [13], researchers explore grounding texts and images to each other to attain multimodal understanding capabilities. Our proposed method also focuses on tuning large models but places a distinct emphasis on robust representation learning.
Fig. 1: The overview of our proposed method. The image and text are first processed by separate encoders for robust representation learning. After that we fuse the modality features and robust embedding for the final predictions.
3 Method
The pipeline of our proposed approach is illustrated in Figure 1. Our method comprises two main modules for image-text classification. Given an image-text pair ($I$, $T$), we initially dispatch them to their corresponding frozen foundation models for feature extraction. In this stage, a modality latent translation module is introduced to facilitate robust representation learning. Subsequently, the obtained robust representation, in conjunction with text and image embeddings, is integrated for final predictions through our newly designed structure.
3.1 Modality Latent Translation
When the image and text are processed by respective large models, a modality latent translation module is introduced to learn the robust representation. For each pair of image-text foundation model layers $l$, the corresponding representations are $I_{l}\in R^{N_{i}\times d_{i}}$ and $T_{l}\in R^{N_{t}\times d_{t}}$, where $N$ and $d$ are respective token numbers and dimensions. Then two linear projections are employed to map the embeddings into the same space:
$I_{l}^{{}^{\prime}}=W_{i}\cdot I_{l}+b_{i}\in R^{N_{i}\times d_{c}},$ (1)
$T_{l}^{{}^{\prime}}=W_{t}\cdot T_{l}+b_{t}\in R^{N_{t}\times d_{c}},$
where $W_{i}\in R^{d_{i}\times d_{c}}$, $W_{t}\in R^{d_{t}\times d_{c}}$, $b_{i}\in R^{d_{c}}$, and $b_{t}\in R^{d_{c}}$ are four learnable parameters, and $d_{c}$ is the dimension of common space. Followingly, two cross-attention module is employed to perform the modality interaction between two embeddings:
$H_{i,l}=SoftMax(\frac{Q_{i}K_{t}^{T}}{\sqrt{d_{c}}})V_{t},$ (2)
$H_{t,l}=SoftMax(\frac{Q_{t}K_{i}^{T}}{\sqrt{d_{c}}})V_{i},$
where $Q_{i}$, $K_{i}$, and $V_{i}$ are transformed modality embeddings from image embedding $I_{l}^{{}^{\prime}}$. The same is true for $Q_{t}$, $K_{t}$, and $V_{t}$. Through cross-attention, the modality embeddings gain access to information from each other, yielding a more comprehensive set of common semantics. The resulting normalized embeddings, augmented with interacted residuals, are then regarded as the representation of each modality in the common space:
$H_{i,l}^{{}^{\prime}}=Norm(H_{i,l}+I_{l}^{{}^{\prime}}),$ (3)
$H_{t,l}^{{}^{\prime}}=Norm(H_{t,l}+T_{l}^{{}^{\prime}}).$
Finally, we try to maximize the canonical correlation between $H_{i,l}^{{}^{\prime}}$ and $H_{t,l}^{{}^{\prime}}$ via DCCA [5], so as to bring them closer to each other in the common space. Specifically, let $R_{11}$, $R_{22}$ be variances of $H_{i,l}^{{}^{\prime}}$ and $H_{t,l}^{{}^{\prime}}$, the covariance between $H_{i,l}^{{}^{\prime}}$ and $H_{t,l}^{{}^{\prime}}$ as $R_{12}$. The canonical-correlation analysis(CCA) loss can be defined by:
$L_{CCA}=-trace(F^{T}F)^{0.5},$ (4)
where $F=R_{11}^{-0.5}R_{12}R_{22}^{-0.5}$. Throughout the tuning process, the representations become more robust as the canonical correlation increases. Consequently, we can effectively translate representations from one modality to another, thereby empowering the model with the ability to infer in scenarios where one modality is missing.
3.2 Fusion and Training Target
After the frozen foundation model, the robust representations ($H_{i,l}^{{}^{\prime}}$ and $H_{t,l}^{{}^{\prime}}$) and extracted modality embeddings($E_{i}$, and $E_{t}$) are sent to our new designed fusion. The detailed structure is shown in Figure 1. We first introduce a learnable vector $M\in R^{L_{s}}$ to average-pool the robust representations from each selected image-text layers:
$H_{l}$ $=\frac{1}{2}(H_{i,l}^{{}^{\prime}}+H_{t,l}^{{}^{\prime}}),$ (5)
$H_{r}$ $=Avg(M\cdot[H_{1},H_{2},...,H_{l},...,H_{L_{s}}]),$
where $H_{l}$ is the joint robust representation of layer $l$ and $L_{s}$ is the number of selected layers for robust representation learning. Then we perform the information exchange between $H_{r}$, $E_{i}$, and $E_{t}$:
$E_{r,i}=SoftMax(\frac{Q_{r}K_{i}^{T}}{\sqrt{d_{c}}})V_{i},$ (6)
$E_{r,t}=SoftMax(\frac{Q_{r}K_{t}^{T}}{\sqrt{d_{c}}})V_{t},$
where $Q_{r}$ is projected by $H_{r}$, $K_{i}$/$V_{i}$ are projected from $E_{i}$, and $K_{t}$/$V_{t}$ are projected from $E_{t}$. Finally, we utilize the mean of both for predictions:
$\hat{y}=Classifier(\frac{1}{2}(E_{r,i}+E_{r,t})).$ (7)
In our method, we employ two training targets: the CCA loss and the task loss. The final loss function can be represented by:
$L=\alpha L_{CCA}+\beta L_{CE},$ (8)
where $L_{CE}$ is the cross entropy loss, $\alpha$ and $\beta$ are two hyper-parameters.
4 Experiments and Analysis
4.1 Benchmark Datasets
To evaluate the effectiveness of our proposed method, we conduct the experiments on three public datasets: MM-IMDB [14], UPMC-Food101 [15], and SNLI-VE [16]. Among the three datasets, MM-IMDB dataset is to classify the movie into one or more of the 23 genres with the poster image and textual outlines. This dataset contains contains 15510 training samples, 2599 validation samples and 7779 samples for test. UPMC-Food101 is a popular image-text classification dataset, which aims to categorize food images with recipe descriptions into 101 categories. There are 67971 training samples and 22715 test samples. SNLI-VE is a visual-entailment understanding dataset, in which each sample includes an image premise and a text hypothesis. The labels are annotated by the semantic relationship(entailment, neutral, or contradiction) between them. The datasets contains 529527 samples for training, 17585 for validation, and 17901 for test.
4.2 Experimental Settings
In our experiments, we employ the pretrained LLaMA as the text foundation model and the image encoder of CLIP-L/224 as visual foundation model. Inherited from the pretrained models, $d_{i}$ and $d_{t}$ are set to 4096 and 768, respectively. The dimension of common space $d_{c}$ is set to 1024. We infuse the MoLT module in the last 4 layers of image and text models, meaning that $L_{s}$ is 4. In the loss function, we set $\alpha$ to 0.1 and $\beta$ to 0.9. To reduce the memory consumption, we train our model with mixed-precision. The Adam optimizer is employed in our method and the learning rate is set to 0.0004. Our approach is implemented with PyTorch framework and the experiments are conducted on two NVIDIA RTX 3090Ti GPUs.
Table 1: The quantitative results of our method on three benchmark datasets. w/ LM indicates whether the large models are utilized in the approach.
Method w/ LM MM-IMDB UPMC-Food101 SNLI-VE
F1-micro/macro(%) Acc(%) Acc(%)
HUSE [17] - 92.30 -
VisualBERT [18] - 92.30 75.06
MMBT [19] $\checkmark$ 66.8 / 61.8 92.10 74.69
MaPLe [20] $\checkmark$ 60.9 / 51.2 90.80 71.52
BlindPrompt [21] $\checkmark$ 56.5 / 50.2 84.56 65.54
PMF [22] $\checkmark$ 64.5 / 58.8 91.51 71.92
Ours $\checkmark$ 64.9/59.0 92.12 75.10
Modality-absence Inference
Baseline $\checkmark$ 59.0 / 51.2 85.2 69.3
Ours(text-absence) $\checkmark$ 62.4 / 56.7 88.9 73.2
Ours(image-absence) $\checkmark$ 63.1 / 57.0 89.2 71.0
Table 2: The ablation study on SNLI-VE dataset. C.A. means the cross-attention in MoLT. Fusion indicates the fusion strategy in our method.
Ablation SNLI-VE
C.A. $L_{CCA}$ $M$ Fusion Acc(%)
$\checkmark$ $\checkmark$ $\checkmark$ 69.6
$\checkmark$ $\checkmark$ $\checkmark$ 70.5
$\checkmark$ $\checkmark$ $\checkmark$ 73.0
$\checkmark$ $\checkmark$ $\checkmark$ 71.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 73.45
Fig. 2: The visualization of some cases for our propose method and the baseline model.
4.3 Quantitative Results
The results of our method on the evaluated datasets are presented in Table 1. As evident from the results, we achieve state-of-the-art performance on each benchmark. Among the methods we compare with, HUSE [17] and VisualBert [18] do not utilize large models, while others (such as MaPLe [20], MMBT [19], and PMF [22]) are based on large models. Most large model-based methods aim to facilitate information exchange through fine-tuning. The performance gap observed among them underscores the potent generalization ability of large models.
4.4 Ablation Study
To further investigate the effectiveness of each component in our method, we conducted an ablation study using the SNLI-VE dataset. The results are presented in Table 2. When removing $L_{CCA}$, the performance drops dramatically, highlighting the crucial role of the training target. The results also demonstrate that the cross-attention module and the learnable vector $M$ have a positive influence on the final outcomes. When the fusion module is not utilized, meaning only the robust representations are employed for predictions, the performance also decreases. The performance gap observed between each ablative model and our final model underscores the effectiveness of our proposed method.
4.5 Robust Inference Analysis
As our focus lies in enhancing model robustness in modality-missing scenarios, we conducted corresponding experiments. Table 1 also presents the performance when inferring with only one modality. The results show that each modality has a varying impact on different tasks. For movie classification, the textual modality dominates, while the importance of the image modality increases for visual-entailment understanding. Nevertheless, the performance is consistently better than the baseline, which does not utilize the MoLT module and does not incorporate robust representation learning, thereby revealing the effectiveness of our proposed approach. Additionally, we visually inspect some cases in Figure 2. With the MoLT module and robust representation learning, the model can still predict results accurately, whereas the baseline model often fails.
5 Conclusion
In this paper, we propose a robust representation learning strategy tailored for large models. Our approach incorporates a modality latent translation module capable of translating one modality embedding to another. Additionally, we introduce a novel fusion schema for robust representation and modality embeddings. The experiments are conducted on three datasets, and the results clearly illustrate the effectiveness of our proposed method. In the future, we plan to conduct further research in robust representation learning to enhance our ability to handle modality-absence scenarios more effectively.